97 students at a college were asked whether they had completed their required English 101 course, and 54 students said "yes". Construct the 80% confidence interval for the proportion of students at the college who have completed their required English 101 course. Enter your answer as a tri-linear inequality using decimals (not percents) accurate to three decimal places.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 54 / 97 = 0.558

1 - \hat p = 0.442

Z\alpha/2 = 1.282

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.282 * \sqrt ((0.558 * 0.442) / 97)

= 0.065

A 80 % confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.558 - 0.065 < p < 0.558 + 0.065

0.493 < p < 0.623

(0.493, 0.623)